please help...Using factoring to solve this quadratic equation:
4x^2-17x-15=0

Answer by venugopalramana(3286)

please help...Using factoring to solve this quadratic equation:
4x^2-17x-15=0
=4x^2-20x+3x-15=0
=4x(x-5)+3(x-5)=0
=(x-5)(4x+3)=0
hence
x-5=0...that is x=5...or
4x+3=0...that is 4x=-3...x = -3/4